A person scores a 81 on a test of verbal ability and a 6.4 on a tesat of quantitative ability. For the verbal ability test, the mean for people in general is 50 and the standard deviation is 20. For the quantitative ability test, the mean for people in general is 0 and the standard deviation is 5. Which is the persons stronger ability: verbal of quatitative? explain your answer to a person who has never hada course in statistics.
Answer by stanbon(75887)

A person scores a 81 on a test of verbal ability and a 6.4 on a test of quantitative ability. For the verbal ability test, the mean for people in general is 50 and the standard deviation is 20. For the quantitative ability test, the mean for people in general is 0 and the standard deviation is 5. Verbal: z(81) = (81-50)/20 = 31/20 = 1.55
----
Quantitative: z(6.4) = (6.4-0)/5 = 1.28
-----
Which is the persons stronger ability: verbal of quatitative?
---
Ans: verbal
=========================
Cheers,
Stan H.
====================